OpenStudy (anonymous):

What is 0.000001 written as a power of ten?

OpenStudy (anonymous):

take a shot on this one.

OpenStudy (the_fizicx99):

1 * 10^-6

OpenStudy (the_fizicx99):

Scientific notation

OpenStudy (anonymous):

i got it its 10^-7

OpenStudy (the_fizicx99):

Why ^-7?

OpenStudy (anonymous):

7 zeros

OpenStudy (anonymous):

\[0.1=\frac{ 1 }{10 }=\frac{ 1 }{ 10^{1} }=10^{-1}\] \[0.01=\frac{ 1 }{ 100 }=\frac{ 1 }{ 10^{2} }=10^{-2}\] \[0.001=\frac{ 1 }{ 1000 }=\frac{ 1 }{ 10^{3} }=10^{-3}\] ................................................................................. \[0.000001=\frac{ 1 }{1000000 }=\frac{ 1 }{ 10^{6} }=10^{-6}\]

OpenStudy (anonymous):

awhwhw

OpenStudy (wolf1728):

freakenpurplepizza You count how many places the decimal has been shifted (see attached).